The firm's fixed cost is given as #1000 and over head coat of #12 per CD produced. Given the demand function for the firm
P=30-0.2√q
Where q is quantity produces and p is price. If the firm aims profit (π) of #20000 per month. Howany quantity (q) should be produced to realize this firm?
1
Expert's answer
2020-05-01T10:22:02-0400

Profit to be achieved set by the firm at # 20,000, fixed costs at # 1000 and additional overhead cost per cd produced at # 12

Firm demand function "p = 30-0.2\sqrt{q}"

"Profit(Pq)=20000=R(q)\u2013C(q)"


Where R (q) is the revenue and C (q) is the total cost P is the price and q is the quantity


"R(q)=q(30−0.2\sqrt{q}=30q–0.2q^\frac{3}{2}"

"C(q)=1000+12q"

Profit equation = "30q–0.2q^\frac{3}{2}" "\u2212(1000+12q)"

"0=18q\u2212""0.2q^\frac{3}{2}" "\u22121000"


Maximum profit can be only found when the first derivative "\int" q= 0

"\int\limits_{x\in C}dx" q="(90q−q^\frac{3}{2}−5000)" "\int\limits_{x\in C}dx"


"0= 90- \frac{3}{2} q^\frac{1}{2}"


"q^\frac{1}{2}" =90"\times \frac{2}{3}"


"q=60^{2}=" 3600

quantity= 3600


This is the maximum quantity that can be produced to realize maximum profit that will be higher that # 20000